# Thinking Models Learn When

### v1

@@ 1 Introduction
Prior work has suggested several hypotheses: (1) thinking models acquire entirely new reasoning capabilities through specialized training (Gandhi et al., 2025); (2) reinforcement learning (RL) teaches them to structure their reasoning more effectively (Marjanović et al., 2025); (3) RL teaches them to repurpose pre-existing base model representations for new mechanisms (Ward et al., 2025); or (4) the additional inference time simply allows more computation to be applied to difficult problems (Zhao et al., 2025; Wang et al., 2025). In this paper, we present evidence for a more nuanced explanation: not only do base models already possess the fundamental reasoning capabilities, but thinking models learn when to deploy these capabilities in a structured sequence.
@@
This finding provides strong evidence that reinforcement learning with verifiable rewards (RLVR (Yue et al., 2025)) used to train thinking models primarily teaches when to activate pre-existing skills rather than teaching how to execute those skills. This perspective has direct implications for more efficient training of reasoning in future language models. To ease reproducibility and further research, we publish our codebase and results in a public GitHub repository [^1] https://github.com/cvenhoff/cot-interp .
@@ 3 Steering Base Models to Reason
In this section, we explore the main question of our paper: do base models already possess the reasoning mechanisms of thinking models, and if so, can we induce these behaviors through targeted interventions? Our hypothesis, supported by preliminary evidence in prior work (Ward et al., 2025; Hou et al., 2023; Galichin et al., 2025), is that non-thinking models may already contain the latent capacity for sophisticated reasoning patterns, such as uncertainty estimation and backtracking, but lack the ability to effectively determine when to employ these mechanisms.
@@ 3.3 Hybrid Model Results
Qwen2.5-14B DeepSeek-R1-Distill-Qwen-14B 59.1% 74.6% (+15.5%) 85.7% (+26.6%) $\mathbf{58.2\%}$
@@
Qwen2.5-14B DeepSeek-R1-Distill-Qwen-14B 6.5% 8.5%
@@ 3.4 Hybrid Model Ablation Studies
The results are concise: Only-bias achieves $76.8\%$, indicating the bias helps but category-specific steering vectors are necessary; Random-firing reaches $77.8\%$, showing that proper timing of activation is crucial; Random-vectors achieves $77.2\%$, confirming that the learned directions are specific rather than generic. Together, these findings support that effectiveness comes from specific learned directions applied with correct, category-timed activation, consistent with our claim that thinking models primarily learn when to deploy reasoning mechanisms.
@@ 4 Related Work
A complementary line asks whether base models contain latent reasoning. Zhao et al. (2025) find RL post-training primarily amplifies pretraining patterns rather than teaching new skills. Wang et al. (2025) show that one carefully chosen example can markedly improve reasoning, suggesting minimal intervention can unlock latent reasoning “circuits”. Minder (2024) investigates how fine-tuning surfaces pre-existing capabilities rather than constructing new mechanisms, identifying shared subspaces between base and fine-tuned models that control context-sensitivity behavior. Venhoff et al. (2025) identify interpretable activation vectors corresponding to reasoning behaviors in thinking models, which can increase or decrease these mechanisms without additional fine-tuning. Ward et al. (2025) discover directions in base models that steer reasoning behavior in thinking LLMs. To our knowledge, no prior work has systematically discovered steering vectors that induce reasoning directly in the base model itself.
On distillation, Baek & Tegmark (2025) identify feature directions in distilled models that steer different thinking modes, and Galichin et al. (2025) use Sparse Autoencoders to interpret reasoning features and enhance capabilities. Concurrently, Bogdan et al. (2025) support sentence-level decomposition of chain-of-thought, showing sentences serve distinct functions. Bridging base and reasoning-specialized models, Jia et al. (2025) integrates a learned latent action space to guide RL fine-tuning, and Zhang et al. (2025) surveys reasoning-centric LLMs, emphasizing high-quality reasoning data and hybrid training.
@@ Acknowledgements
To ensure full reproducibility of our results, we provide comprehensive implementation details and resources. Our complete codebase is available at our GitHub repository [^3] https://github.com/cvenhoff/cot-interp , including: (1) SAE training and evaluation code with exact hyperparameters, layer specifications, and dictionary sizes used for all experiments; (2) Steering vector optimization implementation with training procedures, loss functions, and convergence criteria detailed in Appendix˜C; (3) Complete prompt templates for category generation, consistency evaluation, completeness scoring, and independence assessment used in our taxonomy evaluation pipeline; (4) Dataset processing scripts with exact train/validation splits, random seeds, and preprocessing steps for MMLU-Pro reasoning traces; (5) Hybrid model evaluation framework including the dynamic steering application logic and perplexity-based selection mechanisms; (6) One-command reproduction scripts that reproduce all key experimental results from SAE training through final performance evaluation. All experimental configurations, model checkpoints, and evaluation datasets are documented with version control to ensure exact reproducibility across different computational environments.
@@ C.2 Prompt Template
C.3 Results on Smaller Models Without Sweep
@@
Qwen2.5-14B DeepSeek-R1-Distill-Qwen-14B 59.1% 69.4% (+10.3%) 85.7% (+26.6%) $\mathbf{38.7\%}$
@@ Appendix D Steering Vector Usage Statistics
For the Llama-3.1-8B model with DeepSeek-R1-Distill steering on GSM8K, steering vectors were applied to an average of $125.96$ tokens per problem out of $1641.56$ total tokens, representing $7.7\%$ of the generation process. The most frequently activated reasoning mechanism was “Planning Next Steps” ($18.7\%$ of steered tokens), followed by “Confirming Reasoning Steps” ($16.2\%$) and “Numeric Calculation Steps” ($13.8\%$). The majority of steering applications used a coefficient of $0.3$ ($68.6\%$), indicating that relatively gentle nudges were most effective for this configuration.
@@
On the more challenging MATH500 benchmark, the same model configuration showed different steering patterns. Steering was applied to $91.55$ tokens per problem on average out of $1114.06$ total tokens ($8.2\%$ steered fraction). “Planning Next Steps” remained the most used mechanism ($21.0\%$), and “Mathematical Computation Steps” was equally prominent ($21.0\%$), reflecting the increased mathematical complexity. The steering coefficient distribution shifted toward $0.5$ ($85.2\%$), suggesting that stronger interventions were needed for these harder problems.
@@
The Qwen2.5-14B configuration showed a markedly different pattern, with steering applied to $45.49$ tokens per problem out of $738.91$ total tokens ($6.2\%$ steered fraction). The steering was dominated by “Problem Restatement” ($55.7\%$) and “Metacognitive Markers” ($41.7\%$), with minimal use of computational reasoning mechanisms.
@@
On MATH500, the same model maintained a similar pattern with $62.33$ steered tokens per problem out of $823.96$ total tokens ($7.6\%$ steered fraction). “Problem Restatement” remained dominant ($67.0\%$), followed by “Metacognitive Markers” ($29.8\%$). The steering coefficient remained primarily at $0.5$ ($77.3\%$).
@@
The largest model configuration with DeepSeek-R1-Distill showed relatively low steering intensity, applying vectors to $32.02$ tokens per problem out of $622.16$ total tokens ($5.1\%$ steered fraction). The most frequent mechanism was “Evaluating Expressions” ($29.8\%$), followed by “Drawing Conclusions” ($23.6\%$). The coefficient distribution favored $0.5$ ($67.6\%$).
@@
On MATH500, this configuration steered $54.18$ tokens per problem out of $800.09$ total tokens ($6.8\%$ steered fraction). “Evaluating Expressions” became even more dominant ($43.3\%$), with “Computational step initiation” ($17.5\%$) and “Drawing Conclusions” ($17.2\%$) also playing significant roles. The steering coefficient remained primarily at $0.5$ ($74.4\%$).
@@
The QwQ-32B thinking model produced a different steering pattern, with $30.82$ steered tokens per problem out of $300.96$ total tokens ($10.2\%$ steered fraction). “Numeric computations” was the most frequent mechanism ($28.9\%$), followed by “Stating Known Equations” ($28.2\%$). The coefficient distribution was heavily concentrated at $0.5$ ($88.3\%$).
@@
On the challenging MATH500 benchmark, this configuration showed higher steering intensity with $98.08$ steered tokens per problem out of $920.49$ total tokens ($10.7\%$ steered fraction). The mechanism distribution was balanced, with “Stating Known Equations” ($24.3\%$), “Presenting Conclusions” ($21.3\%$), “Numeric computations” ($20.7\%$), and “Planning Next Steps” ($20.5\%$) all playing substantial roles. The steering coefficient remained primarily at $0.5$ ($69.7\%$), with increased use of higher coefficients ($0.9$: $8.4\%$, $0.8$: $6.8\%$).

### v2

@@ 1 Introduction
Prior work has suggested several hypotheses: (1) thinking models acquire entirely new reasoning capabilities through specialized training (Gandhi et al., 2025); (2) reinforcement learning (RL) teaches them to structure their reasoning more effectively (Marjanović et al., 2025); (3) RL teaches them to repurpose pre-existing base model representations for new mechanisms (Ward et al., 2025a); or (4) the additional inference time simply allows more computation to be applied to difficult problems (Zhao et al., 2025; Wang et al., 2025). In this paper, we present evidence for a more nuanced explanation: not only do base models already possess the fundamental reasoning capabilities, but thinking models learn when to deploy these capabilities in a structured sequence.
@@
This finding provides strong evidence that reinforcement learning with verifiable rewards (RLVR (Yue et al., 2025)) used to train thinking models primarily teaches when to activate pre-existing skills rather than teaching how to execute those skills. This perspective has direct implications for more efficient training of reasoning in future language models. To ease reproducibility and further research, we publish our codebase and results in a public GitHub repository [^1] https://github.com/cvenhoff/thinking-llms-interp .
@@ 3 Steering Base Models to Reason
In this section, we explore the main question of our paper: do base models already possess the reasoning mechanisms of thinking models, and if so, can we induce these behaviors through targeted interventions? Our hypothesis, supported by preliminary evidence in prior work (Ward et al., 2025a; Hou et al., 2023; Galichin et al., 2025), is that non-thinking models may already contain the latent capacity for sophisticated reasoning patterns, such as uncertainty estimation and backtracking, but lack the ability to effectively determine when to employ these mechanisms.
@@ 3.3 Hybrid Model Results
Qwen2.5-14B DeepSeek-R1-Distill-Qwen-14B 58.6% 75.4% (+16.8%) 86.4% (+27.8%) $\mathbf{60.4\%}$
@@
Qwen2.5-14B DeepSeek-R1-Distill-Qwen-14B 6.5% 7.8%
@@ 3.4 Hybrid Model Ablation Studies
As a reminder, the base model Qwen2.5-32B achieves $63.4\%$ accuracy and the full hybrid model achieves $84.4\%$ accuracy on this benchmark. The ablation results show that Only-bias achieves $76.8\%$, indicating the bias helps but category-specific steering vectors are necessary; Random-firing reaches $77.8\%$, showing that proper timing of activation is crucial; Random-vectors achieves $77.2\%$, confirming that the learned directions are specific rather than generic. A further analysis of the behavior induced by the bias vector shows that it captures higher-level stylistic and structural aspects of thinking model outputs, such as providing more self-contained and pedagogical explanations (see Section˜C.3). Together, these findings indicate that the effectiveness of thinking models can only be partially explained by a more verbose answer style, and closing the gap requires specific learned directions applied with correct, category-timed activation, consistent with our claim that thinking models primarily learn when to deploy reasoning mechanisms.
@@ 4 Related Work
[…]
On distillation, Baek & Tegmark (2025) identify feature directions in distilled models that steer different thinking modes, and Galichin et al. (2025) use Sparse Autoencoders to interpret reasoning features and enhance capabilities. Troitskii et al. (2025) analyze how internal features preceding “wait” tokens modulate reasoning patterns in DeepSeek-R1-Distill using sparse crosscoders, finding that features both promoting and suppressing wait tokens influence downstream reasoning behavior. Concurrently, Bogdan et al. (2025) support sentence-level decomposition of chain-of-thought, showing sentences serve distinct functions. Bridging base and reasoning-specialized models, Jia et al. (2025) integrates a learned latent action space to guide RL fine-tuning, and Zhang et al. (2025) surveys reasoning-centric LLMs, emphasizing high-quality reasoning data and hybrid training.
@@ Acknowledgements
To ensure full reproducibility of our results, we provide comprehensive implementation details and resources. Our complete codebase is available at our GitHub repository [^3] https://github.com/cvenhoff/thinking-llms-interp , including: (1) SAE training and evaluation code with exact hyperparameters, layer specifications, and dictionary sizes used for all experiments; (2) Steering vector optimization implementation with training procedures, loss functions, and convergence criteria detailed in Appendix˜C; (3) Complete prompt templates for category generation, consistency evaluation, completeness scoring, and independence assessment used in our taxonomy evaluation pipeline; (4) Dataset processing scripts with exact train/validation splits, random seeds, and preprocessing steps for MMLU-Pro reasoning traces; (5) Hybrid model evaluation framework including the dynamic steering application logic and perplexity-based selection mechanisms; (6) One-command reproduction scripts that reproduce all key experimental results from SAE training through final performance evaluation. All experimental configurations, model checkpoints, and evaluation datasets are documented with version control to ensure exact reproducibility across different computational environments.
@@ C.2 Prompt Template
C.3 Interpreting the Bias Vector
To understand what behaviors the general bias vector captures, we conducted an interpretability experiment comparing outputs from the base model (Qwen2.5-32B) with outputs from a hybrid model using only the bias vector (no category-specific steering). We analyzed $500$ paired responses on the MATH500 benchmark, where each pair consists of the base model’s response and the bias-only hybrid model’s response to the same question.
For each batch of $50$ question-response pairs, we prompted an LLM evaluator (OpenAI’s o4-mini) to identify qualitative differences in writing style, reasoning approach, and structural patterns between the two model variants, explicitly avoiding judgments about numerical accuracy or correctness. The evaluator focused on consistent behavioral differences across batches rather than problem-specific variations. After collecting these batch-level summaries, we synthesized them into a consolidated description of the bias vector’s effect.
The analysis revealed that the bias vector induces a systematic shift in how the base model presents its reasoning. Specifically, across all batches, the base model adopts a sparse, mechanical outline (usually a numbered list of procedural steps with minimal narrative or justification), often jumping straight to formulas or computations and sometimes truncating context or checks. The bias-only hybrid model, in contrast, weaves its solutions into more discursive, conversational expositions: restating the problem, providing high-level overviews, explaining rationale, recalling relevant theorems, and verifying intermediate results. Structurally, the base model favors uniform “Step $1$, Step 22…” bullet points, whereas the bias-only hybrid mixes prose with headings or subheadings to produce a more cohesive tutorial-style write-up. Qualitatively, the base model feels brisk and utilitarian, while the bias-only hybrid reads as a self-contained guided walkthrough with fuller explanations.
These findings suggest that the bias vector captures higher-level stylistic and structural aspects of thinking model outputs (such as using a more verbose narrative, providing contextual framing, and adopting a more pedagogical tone) rather than specific reasoning mechanisms. This is consistent with the bias vector’s intended role as a general scaffold that encourages the base model to produce more elaborate, explanatory reasoning traces that resemble the format of thinking model outputs. The complete experimental procedure and results are available in the supplementary code repository.
C.4 Results on Smaller Models Without Sweep
@@
Qwen2.5-14B DeepSeek-R1-Distill-Qwen-14B 58.6% 69.4% (+10.8%) 86.4% (+27.8%) $\mathbf{38.8\%}$
@@ Appendix D Steering Vector Usage Statistics
For the Llama-3.1-8B model with DeepSeek-R1-Distill steering on GSM8K, steering vectors were applied to an average of $73.58$ tokens per problem out of $523.59$ total tokens, representing $14.1\%$ of the generation process (or $21.5\%$ per-problem average). The most frequently activated reasoning mechanism was “Planning Next Steps” ($17.8\%$ of steered tokens), followed by “Numeric Calculation Steps” ($17.0\%$) and “Task Formulation” ($12.4\%$). The majority of steering applications used a coefficient of $0.1$ ($67.9\%$), indicating that relatively gentle nudges were most effective for this configuration.
@@
On the more challenging MATH500 benchmark, the same model configuration showed different steering patterns. Steering was applied to $91.96$ tokens per problem on average out of $1036.17$ total tokens ($8.9\%$ steered fraction, or $15.8\%$ per-problem average). “Planning Next Steps” remained the most used mechanism ($21.2\%$), and “Mathematical Computation Steps” was equally prominent ($20.5\%$), reflecting the increased mathematical complexity. The steering coefficient distribution shifted toward $0.1$ ($69.7\%$), with additional usage of higher coefficients ($0.2$: $3.9\%$, $0.3$: $3.8\%$, $0.5$: $3.6\%$).
@@
The Qwen2.5-14B configuration showed a markedly different pattern, with steering applied to $82.30$ tokens per problem out of $1637.47$ total tokens ($5.0\%$ steered fraction, or $6.5\%$ per-problem average). The steering was dominated by “Problem Restatement” ($50.3\%$) and “Metacognitive Markers” ($47.1\%$), with minimal use of computational reasoning mechanisms ($2.4\%$ for “Numeric Computation”). The steering coefficient distribution was split between $0.1$ ($48.5\%$) and higher coefficients ($0.8$: $6.4\%$, $0.7$: $6.2\%$, $0.9$: $6.1\%$).
@@
On MATH500, the same model maintained a similar pattern with $85.33$ steered tokens per problem out of $1445.58$ total tokens ($5.9\%$ steered fraction, or $7.8\%$ per-problem average). “Problem Restatement” remained dominant ($60.7\%$), followed by “Metacognitive Markers” ($35.5\%$) and “Numeric Computation” ($3.8\%$). The steering coefficient distribution favored $0.1$ ($39.5\%$) and $0.5$ ($22.1\%$), with significant usage of higher coefficients ($0.9$: $5.8\%$, $0.8$: $5.7\%$, $0.7$: $5.2\%$, $1.0$: $5.1\%$).
@@
The largest model configuration with DeepSeek-R1-Distill showed relatively low steering intensity, applying vectors to $31.92$ tokens per problem out of $603.52$ total tokens ($5.3\%$ steered fraction, or $11.2\%$ per-problem average). The most frequent mechanism was “Evaluating Expressions” ($30.8\%$), followed by “Drawing Conclusions” ($23.5\%$) and “Stating Given Problem Data” ($11.7\%$). The coefficient distribution heavily favored $0.5$ ($68.4\%$), with additional usage of higher coefficients ($0.7$: $7.0\%$, $0.6$: $6.6\%$, $0.8$: $6.4\%$).
@@
On MATH500, this configuration steered $54.23$ tokens per problem out of $783.54$ total tokens ($6.9\%$ steered fraction, or $9.8\%$ per-problem average). “Evaluating Expressions” became even more dominant ($43.0\%$), with “Computational step initiation” ($17.5\%$) and “Drawing Conclusions” ($16.6\%$) also playing significant roles. The steering coefficient remained primarily at $0.5$ ($74.4\%$), with additional usage of higher coefficients ($0.7$: $5.7\%$, $0.6$: $5.3\%$, $0.8$: $5.1\%$, $0.9$: $5.0\%$).
@@
The QwQ-32B thinking model produced a different steering pattern, with $30.85$ steered tokens per problem out of $300.37$ total tokens ($10.3\%$ steered fraction, or $12.7\%$ per-problem average). “Numeric computations” was the most frequent mechanism ($28.8\%$), followed by “Stating Known Equations” ($28.2\%$) and “Presenting Conclusions” ($17.9\%$). The coefficient distribution was heavily concentrated at $0.5$ ($88.2\%$), with minimal usage of higher coefficients ($1.0$: $2.8\%$, $0.6$: $2.4\%$).
@@
On the challenging MATH500 benchmark, this configuration showed higher steering intensity with $98.08$ steered tokens per problem out of $920.49$ total tokens ($10.7\%$ steered fraction, or $12.0\%$ per-problem average). The mechanism distribution was balanced, with “Stating Known Equations” ($24.3\%$), “Presenting Conclusions” ($21.3\%$), “Numeric computations” ($20.7\%$), and “Planning Next Steps” ($20.5\%$) all playing substantial roles. The steering coefficient remained primarily at $0.5$ ($69.7\%$), with increased use of higher coefficients ($0.9$: $8.4\%$, $0.8$: $6.8\%$, $1.0$: $6.7\%$).
D.9 Qwen2.5-Math-1.5B with DeepSeek-R1-Distill on GSM8K
For the specialized Qwen2.5-Math-1.5B model with DeepSeek-R1-Distill steering on GSM8K, steering vectors were applied to an average of $42.83$ tokens per problem out of $464.49$ total tokens, representing $9.2\%$ of the generation process (or $12.8\%$ per-problem average). The most frequently activated reasoning mechanism was “Planning Next Step” ($20.4\%$ of steered tokens), followed by “Announcing Intermediate Conclusions” ($16.4\%$) and “Intermediate Numeric Calculations” ($15.8\%$). The majority of steering applications used a coefficient of $0.1$ ($51.9\%$), with fairly distributed usage across other coefficients ($0.3$: $6.0\%$, $0.5$: $6.0\%$, $0.2$: $5.9\%$).
D.10 Qwen2.5-Math-1.5B with DeepSeek-R1-Distill on MATH500
On MATH500, the specialized model showed increased steering with $64.16$ tokens per problem out of $757.04$ total tokens ($8.5\%$ steered fraction, or $10.7\%$ per-problem average). “Planning Next Step” remained the most frequent mechanism ($25.3\%$), followed by “Task Formulation” ($23.1\%$) and “Announcing Intermediate Conclusions” ($13.4\%$). The steering coefficient distribution remained primarily at $0.1$ ($53.2\%$), with distributed usage across other coefficients ($0.2$: $6.0\%$, $0.4$: $6.0\%$, $0.5$: $5.8\%$).